OpenStudy (ashleyisakitty):

If a vector has an initial point of (2,6) and a terminal point (14,11). Out of 5, 11, 12, and 13, what is closest to the magnitude of the vector?

OpenStudy (e.mccormick):

Know how to find the magnitude?

OpenStudy (anonymous):

Use the distance formula.

OpenStudy (anonymous):

The distance between the points will give the magnitude of the vector.

OpenStudy (e.mccormick):

Distance formula, aka Pythagorean Theory applied to distance: \(D=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) The magnitude of a vector is equal to the distance it travels.

OpenStudy (e.mccormick):

Well, that was some interesting lag.

OpenStudy (ashleyisakitty):

I did the distance formula, and I believe the answer is 12. Is that true?

OpenStudy (e.mccormick):

\(D=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) \(D=\sqrt{(14-2)^2+(11-6)^2}\) \(D=\sqrt{(12)^2+(5)^2}\) \(D=\sqrt{144+25}\) \(D=\sqrt{169}\) Find your mistake?